The half life of iridium is 12.8 years. Calculate decay constant and average life in minutes


1
Expert's answer
2022-05-26T12:07:03-0400

Solution:

iridium (Ir)

The half-life (t1/2) of iridium is 12.8 years


Convert years to minutes:

minutes = years × 525949.2

Therefore,

(12.8 years) × (525949.2) = 6732149.76 min = 6.732×106 min

The half-life (t1/2) of iridium is 6.732×106 min


From law of radioactive decay:

t1/2 = ln(2) / λ

where λ = decay constant

λ = ln(2) / t1/2

λ = (0.693) / (6.732×106 min) = 10.3×10−8 min−1

λ = 10.3×10−8 min−1


Tλ = 1

where T = average life

T = 1 / λ

T = 1 / (10.3×10−8 min−1) = 9.7×106 min

T = 9.7×106 min


Answer:

The decay constant (λ) of iridium is 10.3×10−8 min−1

The average life (T) of iridium is 9.7×106 min